Please can someone solve this problem for me because I tried for a long time to solve it but I couldn't, so the math problem is:
a) use slopes to determine weather each set of points is collinear or non-collinear.
b) if the points are collinear, find the slope.
16) A(-2,1), B(-1,2),C(-4,-2)
17) M(-1,-3),N(2,5),P(3,4)
18) W(-3,-1),X(9,8),Z(5,5)
Collinear means in the same line

Found 2 solutions by ewatrrr, tajar:
Answer by ewatrrr(24785)

16) non-collinear
A(-2,1),
B(-1,2) m = -1/-1 = 1
C(-4,-2) m = 4/3
17) non-collinear
M(-1,-3),
N(2,5) m = -8/-3
P(3,4) m = 1/-1
18) collinear
W(-3,-1),
X(9,8) m = -9/-12 = 3/4
Z(5,5) m = 3/4
W(-3,-1) m = 6/8 = 3/4

Answer by tajar(5)

FIRST RULE to use-Draw the points on coordinate systems to determine if any of the equations are clearly not linear (USE LOGIC)